OpenStudy (anonymous):

Under water, pressure increases 4.3 pounds per square inch for every 10 feet you descend. This represented by the equation p = 0.43d + 14.7, where p is the psi and d is the depth in ft. Find the pressure at 200 feet.

OpenStudy (anonymous):

P=0.43 d +14. 7 P=0.43 200 +14.7 P=83 +14.7 P=100.7 so the pressure at 200 feet is 100.7 psi